OpenStudy (anonymous):

it has been reported that 83% of federal government employees use email. if a sample of 200 federal government employees is selected, find the mean, variance, and standard deviation of the number who use email

OpenStudy (amistre64):

if we can use binomial probability then: mean = np sd = sqrt(npq) and v = npq

OpenStudy (amistre64):

n = 200 p = .83 q = (1-p) = .17

OpenStudy (amistre64):

mean = 166 var = 28.22 sd = 5.3122....

OpenStudy (anonymous):

thank but how did u get that